Calculus II Spring 2002 MA114:004–006 Russell Brown

• Homework B. Due Wednesday, 23 January 2002. Problem #1 on page 433 of the text. Time will be provided in recitation on Tuesday, 22 January 2002 to work on this problem.

• Quiz #2. Thursday, 24 January 2002. Topic: §6.2-4. Basic properties of exponential and logarithm functions. Please be prepared to answer questions similar to the routine homework problems assigned from these sections.

• Homework #4 §6.5 #1, 3, 5, 7, 9, 11, 17. §6.6 #15, 17, 19, 21, 27, 29, 31, 33, 35, 63, 83, 85, 87, 93.

Section 6.5 will be discussed in lecture on Wednesday, 23 January. Be prepared to begin discussing these problems in recitation on Thursday, 24 January. Section 6.6 will be discussed in lecture on Friday, 25 January and in recitation on Tuesday, 29 January.

• Warning. We may continue on to section 7.1 before the first test. Details to follow.

• The notebook of solutions and handouts is finally in the math library in the basement of POT.

• Assignments are available on the web from a link at http://www.math.uky.edu/∼rbrown/courses

January 16, 2002